The first new wells are being drilled in the Ashland Field and the initial new well is completed and will be reported in the very near future.  Obviously Aethon thinks it will be a good well since the company is drilling the second well and holds a permit for a third.  Here is a cut and paste from SONRIS covering all three.  Our members with drilling experience will need to decipher all the driller report abbreviations. The hyperlinks are functional for those that wish to see additional detail.  The

HA RA SUL;WEYERHAEUSER 4-9 HC permit is for six months and expires 6/25.

SERIAL

WELL NAME

WELL NUM

ORG ID

FIELD

PARISH

PROD TYPE

SEC

TWN

RGE

EFFECTIVE DATE

API NUM

252834

HA RA SUE;LA MINERALS 18-19 HC

001

A1760

0131

35

00

018

13N

07W

09/30/2021

17069205000000

 

PRMT DATE

SPUD DATE

STAT DATE

ST CD

06/14/2021

07/09/2021

09/30/2021

31

SCOUT INFO

REPORT DATE

WELL STATUS

MEASURED DEPTH

TRUE VERT DEPTH

DETAIL

11/29/2021

09

20744

13208

SCOUT CODE 8: 11/23/2021 TUES. 6:00 AM - DRILL PLUGS 59 THRU 63. CBU. CHANGE RAMS IN RASU. DRILL PLUGS 64 THRU 81. CBU. ; 11/24/2021 WED. 6:00 AM - FINISH CIRCULATIONG 2 BU. POOH BACK REAMING 14 JTS OF 2 7/8" TBG. CIRCULATE DUE TO CASING PRESSURE CLIMBING TO 7,000 PSI. POOH 2 JTS. REPAIR EQUALIZING VALVE ON RASU. SWAP OUT RAMS. POOH LD 350 JTS.; 11/25/2021 THURS. 6:00 AM - CPOOH. SNUB OUT LD W/ TOTAL OF 653 JTS AND BHA.CLOSE FRAC VALVE, RDMO RASU. ND BOP STACK NU FLOW TREE. RDMO WOR & EQUIPMENT. SECURE WELL, TTS ON FRIDAY.; 11/26/2021 FRI. 6:00 AM - NO ACTIVITY; 11/27/2021 SAT. 6:00 AM - NO ACTIVITY; 11/28/2021 SUN. 6:00 AM - NO ACTIVITY; 11/29/2021 MON. 6:00 AM - NO ACTIVITY

 

 

Well Ser

Well Name

Well Num

Status

Class

Class Type

API Number

Org ID

Permit Date

Sect

Tshp

Rng

Parish

Mer

Lease Num

Total Depth

252834

HA RA SUE;LA MINERALS 18-19 HC

001

31

17069205000000

A1760

06/14/2021

018

13N

07W

35

W

20744

LUW Association

Effective Date

End Date

LUW Code

LUW Type

LUW Name

 

Well Ser

Well Name

Well Num

Status

Class

Class Type

API Number

Org ID

Permit Date

Sect

Tshp

Rng

Parish

Mer

Lease Num

Total Depth

253123

HA RA SUN;ROWELL 16-21 HC

001

01

17069205090000

A1760

12/06/2021

016

13N

07W

35

W

23144

LUW Association

Effective Date

End Date

LUW Code

LUW Type

LUW Name

 

Well Ser

Well Name

Well Num

Status

Class

Class Type

API Number

Org ID

Permit Date

Sect

Tshp

Rng

Parish

Mer

Lease Num

Total Depth

253174

HA RA SUL;WEYERHAEUSER 4-9 HC

001

01

17013222600000

A1760

12/28/2021

033

14N

07W

07

W

21249

LUW Association

Effective Date

End Date

LUW Code

LUW Type

LUW Name

Views: 292

Replies to This Discussion

Explanation of scout report text courtesy of Rock Man.  That TTS near the end is Turn To Sales on Friday.

Pretty simple once you break it down
Obviously tied to drilling out the frac plugs and getting the hole cleaned and ready to flow back post frac
Prior to 11/23, they had already drilled out frac plugs #1 thru 58. Using some sort of drilling assembly on 2.875" tubing
11/23 - first drilled 5 frac plugs / changed out rams (probably due to pressure increase / needed good ram integrity) and then drilled 18 more plugs. Plug 81 appears to be the last frac plug in the lateral / then CBU = circulate bottoms up (hole / lateral is now open to TD 
11/24- continued circulating to clean out hole. Started pulling out of the hole with the tubing and frac plug drill out assembly (circulating as they go to clean out the hole / lateral). Figure there is frac plug debris, frac sand and other material in the hole
Pressured climbed to 7000# during this operation / figure this is flowing pressure since no evidence well was shut in during this operation
Changed out rams again (probably to optimize integrity of rams if they needed to use them)
11/25 - continued to pull of the hole with the tubing and bottom hole frac plug drill out assembly. Laid down all these tubulars. Closed frac valve to "shut in" well.
Rigged down the service rig and moved it off location
Nipple'd down the BOP and replaced it with wellhead tree / Now ready to flow the well back once it is hooked up to sales line or separator / test unit

Aethon spud the Rowell 16-21 HC #2 on January 27.

SONRIS Well Information for serial # 252834 HA RA SUE; LA MINERALS 18-19 HC 001 section 18 13N 07W Natchitoches Parish is showing production info:

04/12/2022 10 20744 13208 COMPLETED 11/26/21; GAS; 11985 MCFD; 18/64 CHOKE; 9225# FP & CP; PERFS 10360-20586'

Thanks, Daniel.  We'll look forward to some monthly production reports.  And the well cost report.

Good looking well on an IP basis.

Surprisingly low water rate (6 BWPD) on IP / would have thought a lot more due to post frac fluid flowback volumes

Production decline now the post important aspect moving ahead

Thanks for posting

RSS

© 2022   Created by Keith Mauck (Site Publisher).   Powered by

Badges  |  Report an Issue  |  Terms of Service